OpenStudy (anonymous):

Hey! :) question What is the length, in feet, of a rectangle hat has a perimeter of 24 feet and a width of 4 feet A) 4 feet B) 8 feet C)12 feet D)32 feet ?

OpenStudy (anonymous):

Let "L" represent length. So Perimeter P = 2(4) + 2L = 8 + 2L = 24. 2L = 16 L = 8 ft

OpenStudy (anonymous):

ohhh cool thanks so much :)